Operations Report – April 24th

Crew 193 Operations Report 24Apr2018

SOL: 2

Name of person filing report: David Attig

Non-nominal systems: One piece EVA suit #10 and two piece EVA suit # 6 have issues

Notes on non-nominal systems: One two piece suit (6) has straps broken holding on neck ring, Visor on suit 10 is heavily scratched and deemed unuseable. See Sol Summary for more details.

Generator (hours run): 12

Solar— SOC 86 % (Before generator is run at night)

Diesel – 75 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0.75 gallons.

Water (trailer) – 0 gallons.

Water (storage) – 550 gallons

Water (static) – 410 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134008.9

Toilet tank emptied: yes
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3):
Honda Used?: Yes
300 Used? Yes
350.1 Used? No
350.2 Used? Yes
350.3 Used? No

ATV Oil Checked?: Yes

ATV Oil Added?: No

ATV Fuel Used: less than 1 Gallon

# Hours the ATVs were Used today: 0.5

Notes on ATVs: With 350.2, throttle cable gets pulled when handlebar is moved all the way to the left. Suggest not using this ATV for future (provided sufficient fuel in others).

Deimos rover used: Yes
Hours: 124.0
Beginning charge: 99%
Ending charge: 98 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 33.7
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 23
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments:

Summary of internet: By 8AM again today, we have used half of our allotted daily internet. Is there a way to monitor our internet over time without constantly logging on to check it? That way we could monitor usage and know if there is a peak time that it is being used.

Summary of suits and radios: Older style (silver front) radios appear to not transmit well enough to use in EVA. One two piece suit broke during EVA today, and 1 one piece suit was deemed unuseable due to scratches in the visor.

Summary of Hab operations: See Sol Summary report.

Summary of GreenHab operations: See GreenHab report

Summary of ScienceDome operations: See Science report.

Summary of RAM operations: We cut a board to use for a new sign in front of MDRS.

Summary of health and safety issues: None to report.

Questions, concerns and requests to Mission Support: See Sol Summary.

EVA 5 Report – April 24th

Sol 2 Crew 193 EVA 5 Summary

Earth-Date April 24 2018

EVA Gold Crew Members – Gold00 (CO Patil-Sabale) / Gold03 (CE Attig / Gold05 (CGA Shear)

EVA Objectives – Summary

Exploration of Tank Wash – Accomplished

Scouting and Mapping of Emergency Shelter Sites – Mapped

EVA Comfort, Well-Being, Safety and Optimization Study

Sign Material Preparation – Accomplished

Issues

One Beartooth device failed before this EVA, so the (deaf) CGA was forced to take his writable LCD tablet along for communication. The stylus went missing at the beginning of the EVA. For the duration of this EVA, we made do with hand gestures and the note app on our phones to communicate with the CGA. Fortunately, his stylus was recovered at the end of this EVA. We will continue to troubleshoot the Beartooth issue.

ATVs Used

Deimos rover

Blue ATVs #300 and 350.2

Refueled and recharged the vehicles at EVA end

Safety Measures & Preparation

Sunscreen/Sun protection – Yes

Adequate hydration – Yes

Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

EVA Objectives – Details

Exploration of Tank Wash – Accomplished

Preparation of Sign Materials in RAM – Accomplished

Emergency Shelter Discovery – Accomplished (at UTM coordinates 0518913, 4252915)

Commander's Report – April 24th

Commander, PHEnOM Gold Crew, MDRS #193

Mission Sol 2 (Earth Day April 24th, 2018)

Woke up to a beautiful Martian morning! Weather on Mars has been beautiful boosting the crew’s exploratory spirits!

After a quick breakfast our first EVA crew, our XO and HSO set on to explore the southern sites of Phobos Peak and Kissing Camel ridge. The crew that stayed back worked on engineering (CE), astronomy (CGA) and greenhab (CO). The crew enjoyed a scrumptious lunch of Thai green vegetable curry and chicken curry with brown rice. It was refreshing to have fresh green bell peppers and tomatoes from our greenhab while cooking those curries!

The afternoon EVA was adventurous as we, the CE, CGA and me, set out to first locate the shelter the previous crew had marked.
We then headed to the Tank Wash area where amongst small canyons, river beds and stream beds we found 2 caves big enough to house 7 to 8 people during an emergency evacuation or to take shelter during storms.

The coordinates for the cave that we think is an ideal emergency shelter are (0518913, 4252915).
We worked in the RAM for the rest of our EVA on the new MDRS signpost project.

The challenge in this EVA was the lack of means to communicate with our CGA who is hearing impaired. His beartooth device malfunctioned. We got creative and used sign language plus typing text in my phone to communicate with him. I feel content knowing we have been able to work around issues, involve him in our EVA’s and he is having as much adventure and fun while working on research and exploration during the EVA’s.

Our CE and I spent some time planning the new MDRS signpost project, we plan to do it in stages.
We are making good progress on our science, gathering useful data. The crew is healthy and doing great!

It has been amazing to see how the crew has been using their skills and expertise to work around problems, come up with unique solutions and make the best of every situation and available tools while working together as a team. Be it fixing and putting up the signpost during an EVA yesterday, or the emergency engineering EVA to secure the RAM for which our HSO volunteered to go in my place or the way she has been using the medical tools available to collect our health vitals before and after the EVA’s, or the way our CE found the GPS unit in the hab and got it working for our GPS Route Measurements project. I am proud to be a Commander of this stellar team!

It is only Sol 2 and I feel happy with the progress we are making already! Will keep the momentum going 🙂

Thank you for being with us Mission Control (MDRS Director) and Mission Support! Appreciate it.

Signing off for this sol,

Warm Regards,
Commander, (Gold 00), Crew #193
Mars Desert Research Station

(P.S. Abbreviations Legend: CO – Commander / XO – Executive Officer / CE – Crew Engineer / HSO – Health Science Officer / CGA – Crew Geologist and Astronomer)

Science Report – April 24th

Here is the Science Report for 24 April 2018

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Both EVAs today had objectives of identifying more sites which would be possible shelters during an emergency evacuation. EVA 4 was able to identify a cave on the south side of west kissing camels ridge which would be an excellent shelter for all possible issues. EVA 5 located the shelter site that was proposed on EVA 1 on the north ridge. EVA 5 also located another possible evacuation site in the Tank Wash area. Further investigation of possible locations will be done on future EVAs.

2. Sunspot and Solar Flare Monitoring: Please refer to the astronomy report.

3. Shortwave Texting on EVA’s: Beartooth devices did not work today. Will need to troubleshoot the devices this evening. Hand sign language was used instead to communicate with Eric as backup means of communication was left on an ATV.

4. GPS Route Measurements: GPS was taken on EVA 5 which has just returned to the hab. Data will be analyzed for a more fulsome report tomorrow.

5. Waterless dish cleaning: No progress today. Initial testing of dirty dish bacteria growth will commence tomorrow with a plan to do testing on every second day due to the amount of supplies brought by gold XO for the research.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working very hard (normally climbing a large hill). Gold 01 is able to track his heart rate over the entire course of the EVA thanks to mobile technology. Results can be exported and used in this study. Analysis will be done as the mission continues.

Sol 2 Summary – April 24th

Crew 193 Sol Summary Report 24APR2018

Sol 2

Summary Title: Every Sol We Hustlin’…

Author: Dr. Shawna Pandya, Crew Health and Safety Officer

Mission Status: Nominal

Sol Activity Summary: EVAs 4 & 5 complete, focusing on science, exploration, and sign-work. In-hab activities consisted of usual maintenance, operations and science. As a team, we managed to fit in some bonding and learn some basic ASL (American Sign Language) phrases from our deaf crewmate. Crucially, I learned to say, “I NEED MY COFFEE!!” in American Sign Language.

Look Ahead Plan: EVA 6 will plan for refueling the ATVs, then North-ward exploration along ‘The Moon’ – objectives: exploration; visor fogging study; emergency shelter scouting on EVA study; crew comfort, health, well-being and safety on EVA study; GPS accuracy on EVA. EVA 7 will plan on continued work in the RAM on the new MDRS sign, as well as visor fogging and crew comfort studies. In-hab activities will focus on the usual maintenance, operations and science. Time permitting, we are for some more crew training in ASL.

Anomalies in work:

1)Team on EVA 5 was forced to use 1 Rover, as the Beartooth communication with our deaf crewmate failed, and the LCD board was taken instead, and could not be safely transported in the ATV.

2)Will be decommissioning spacesuit 7, as the visor is deemed too scratched to allow for safe visualization.

3)One of the two-piece spacesuits came apart at the mechanical neck-ring/backpack interface while on EVA. It will be decommissioned.

Weather: Clear skies in AM, partly cloudy and gusty in PM.

Crew Physical Status: Some blisters, otherwise nominal.

EVA:

EVA 4 – Southeast exploration along Phobos Peak and Kissing Camels’ Ridge E & W; continued work on scouting, crew comfort, visor fogging studies.

EVA 5 – Measurements taken for second MDRS sign; East-ward exploration along Tank Road; continued work on scouting, crew comfort, visor fogging studies.

Reports filed:

GreenHab
EVA 4 Report
Science Report
Astronomy Report

Reports to be filed:
Commander or Journalist’s Report
Operations Report
EVA 5 Report
Photo & Photo of the Day Report

Support Requested:

-Additional fuel for ATVs.

-Requesting location of dropship of supplies that include Vicks Lozenges.

-Happy to attempt to replace the visor on spacesuit 7 if Mission Support can provide us with the appropriate procedure and supplies.

-Happy to attempt to repair the two-piece spacesuit if Mission Support can provide us with the appropriate procedure and supplies.

GreenHab Report – April 24th

Green Hab Report

Name: Doug Campbell

Date: 24 Apr 2018

Environmental control:

Ambient with door open but only for few hours in the afternoon

Shade cloth: on

Fan: not used today

Average temperatures: 39.2 C

Low: 16.7℃

High: 48.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 6.5 Gallons

Time(s) of watering for crops: Twice: In the morning at 900am and afternoon 445pm.

Morning research observations: None

Changes to research plants: An apple seed that sprouted in the hab was transferred to a pot in the green hab.

Aquaponics: None

Narrative: No significant changes.

Support/supplies needed: None

Astronomy Report – April 24th

Astronomy Report

Name: Eric Shear

Crew: 193
Date: April 24, 2018

Sky Conditions: Clear/Cloudy

Wind Conditions: None

Observation Start Time: 0800

Observation End Time: 930

Summary:

Monitored the sun for sunspot activity. Only one sunspot was found – it appears to be a merger of the two seen on the previous observation. Prominences were very small and only visible with the eye, not with the Skyris camera. No flares or filaments were seen.

Photos were taken with the Skyris camera and the best image processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The result is attached.

Objects Viewed: Sun

Problems Encountered: Clouds later in the observation period, after sunspot image was taken. The Registrax program froze and had to be restarted.

EVA Report 1 – April 24th

Sol 2 Crew 193 EVA 4 Summary

Earth-Date April 24 2018

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold004 (HSO Pandya) (Informally: “Team Canada”)

EVA Objectives – Summary

Exploration of Phobos Peak, KIssing Camel Ridge East & West – Accomplished

Scouting and Mapping of Emergency Shelter Sites – Mapped, see below

EVA Comfort, Well-Being, Safety and Optimization Study

Visor Fogging Study – Data collected

Issues

None

ATVs Used

Manual gear-shift (red)

Blue ATV #2

All blue ATVs require fuel; planned for refueling at end of EVA, ran out of time. Will refuel before next use.

Safety Measures & Preparation

Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided

Adequate hydration – Yes

Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

EVA Objectives – Details

Exploration of Phobos Peak, KIssing Camel Ridge East & West – Accomplished

Distance covered by foot – 9.1km

Steps – 12,000

(2) Scouting and Mapping of Emergency Shelter Sites:
Sol: 2
Time of Day: AM
Shelter Designation: Phobos’ Peak & Kissing Camels Ridge E & W
Crew Members: G001/004
Accessibility: Multiple appropriate areas along entire Peak & Ridge
Location: 4250000N, 519000E
Suitability:
Wind storm: Appropriate – Tread with caution, large rocks/potential for landslide

Radiation event: Large, dense, thick rock/caves – seems appropriate, would need to measure depth of rock/caves

Incapacitated crew member: Appropriate

Photo-documentation: Yes
Next Crew Able to Find Shelters based on description/ markers/photodocumentation? N/A

(3) EVA Comfort, Well-Being, Safety and Optimization Study (See below)
Sol 2
Crew Member: G001 | G004
Time of Day: AM | AM
Pre-EVA Weight: 173.4 | 121.2
Post-EVA Weight: 170.4 | 120.6

Pre-EVA BP: 112/76 | 98/70

Post-EVA BP: 120/82 | 106/76
Pre-EVA SpO2: 97% | 96%
Post-EVA SpO2: 93% | 97%

Pre-EVA HR: 67 | 77
Post-EVA HR: 93 | 66
Max Perceived Physical Exertion (1-10): 5 | 5

Max Perceived Mental Exertion (1 (low effort)-10(max concentration)): 1 | 3

Max Perceived Emotional Distress (1 (no distress) – 10(very distressed)): 3 | 2
HR@Max Perceived Exertion: 126 | 136
SpO2 @ Max Perceived Exertion: N/A | 96%
Adequate Pre-Hydration: No | No
Adequate Sun-Protection: Yes | Yes
Subjective Comments: GPS would be really useful | Sunglasses fogged up

(4) Visor Fogging Study – Data Collected

Parameters Choose an option XO HSO EVA Time Morning/Afternoon? AM AM Weather Complete/Partial Overcast Or Clear Clear Clear EVA Suit type One/two piece 1 1 Did the visor fog? Yes/no N N Did the visor fog clear? Yes/no N/A N/A How much did the fog limit visibility? 1-10; clear-opaque N/A N/A What were you doing when it fogged? N/A N/A Did it fog more quickly with exertion? Yes/no N/A N/A What was your exertion level? 1-10; blinking eyes-heart attack N/A N/A How long was the visor fogged? Minutes N/A N/A Did you feel air flow within helmet? Yes/no Rarely Y

EVA Report – April 23rd

Sol 1 Crew 193
EVA 1 Summary
Earth-Date April 23 2018

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold004 (HSO Pandya) (Informally: “Team Canada”)

EVA Objectives – Summary

  1. Marble RItual – Accomplished
  2. Scouting and Mapping of Emergency Shelter Sites – Mapped, see below / Raw data here
  3. Exploration – Accomplished (See below)
  4. EVA Comfort, Well-Being, Safety and Optimization Study / Raw data here
  5. Visor Fogging Study – Data collected / Raw data here
  6. MAG Study – Not collected.

ATVs Used

  • Manual gear-shift (red)
  • Blue ATV #1
  • No issues noted

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – Yes
  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

EVA Objectives – Details

  1. Marble Ritual – Accomplished, no unforeseen events.
  2. Scouting and Mapping of Emergency Shelter Sites:
Omar’s Haven The Rock Slides
Accessibility
  • Far (see map below)
  • Requires moderate level climbing
  • At Marble’s Ritual
  • Immediately accessible
Location
  • See map below (lower blue dot)
  • 4251500N, 518750E
  • See map below (top blue dot)
  • 4250600N, 518700E
Suitability
  • Wind storm: Appropriate
  • Radiation event: RTB suggested
  • Incapacitated crew member: Appropriate

Wind storm: Appropriate Radiation event: RTB suggested Incapacitated crew member: Appropriate PhotodocumentationYY Next crew able to find shelters based on description/markers/photodocumentation

  1. Exploration:
  • Distance covered by foot – 6km
  • Steps – 8,000

hpCNXqiy2AOasKu1dQyVLfKY3goZwVZy8y3L8lE0hpsEPcSUz1Mg15_Xrxh7wJMETcXMMLOJAttK5OR0QGCD146UA40w23NFoDF9h-9yavMqEekl3XyRSS9936yb4SGfK6sf-9jX

  1. EVA Comfort, Well-Being, Safety and Optimization Study (See below)
G001 G004
Time of Day AM AM
Pre-EVA Weight N/A N/A
Post-EVA Weight 172.4lbs 126.4lbs
Pre-EVA BP N/A N/A
Post-EVA BP N/A N/A
Pre-EVA SpO2 N/A N/A
Post-EVA SpO2 N/A N/A
Pre-EVA HR N/A N/A
Post-EVA HR N/A N/A
Max Perceived Physical Exertion (1-10) 6-7 8
Max Perceived Mental Exertion (1 (low effort)-10 (max concentration)) 7-8 3
Max Perceived Emotional Distress (1 (no distress)-10 (very distressed)) 1 1
HR @ Max Perceived Exertion 117 108
Max HR 126 N/A
Sp02 at Max Perceived Exertion N/A 91%

Subjective Improvements for Future EVAs

  • Secure ear-pieces, long hair and sun-protection better
  • Pre-EVA blister prevention suggested prior to next EVA

Sol 1 Crew 193
EVA 2 Summary
Earth-Date April 23 2018

EVA Gold Crew Members – Gold00 (CO Patil-Sabale) / Gold03 (CE Attig)/ Gold05 (CGA Shear)

EVA Objectives – Summary

  1. Marble Ritual – Accomplished
  2. Scouting and Mapping of Emergency Shelter Sites – Mapped, see below
  3. Exploration – Accomplished (See below)
  4. EVA Comfort, Well-Being, Safety and Optimization Study
  5. Emergency rescue of stranded astronomer
  6. Repaired and set up the road sign with existing tools

ATVs Used

  • Blue ATV #1, 2, 3
  • No issues noted

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes
  • Adequate hydration – Yes
  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

EVA Objectives – Details

  1. Marble Ritual – Accomplished, no unforeseen events.
  2. Scouting and Mapping of Emergency Shelter Sites:
Mountain Goat Trail Galileo Road
Accessibility
  • Not immediately accessible, some walking from Miracle Ritual
  • Not immediately accessible, takes some driving
Location
  • 4251000N, 520000E
  • 4252000N, 519000E
Suitability
  • Wind storm: Appropriate
  • Radiation event: RTB suggested
  • Incapacitated crew member: Appropriate
· Wind storm: Appropriate
· Radiation event: RTB suggested
· Incapacitated crew member: Appropriate
Photodocumentation Y Y
Next crew able to find shelters based on description/markers/photodocumentation

3. Exploration:

  • Distance covered by foot – 6.4 km
  • Steps – 8,000

4. EVA Comfort, Well-Being, Safety and Optimization Study (See below)

G00 G03 G05
Time of Day PM PM PM
Pre-EVA Weight 140.2 lbs 222.6 lbs 145.6 lbs
Post-EVA Weight 139.4 lbs 219.4 lbs 144.2 lbs
Pre-EVA BP 126/98 162/110 120/82
Post-EVA BP 118/80 128/88 120/100
Pre-EVA SpO2 97% 93% 95%
Post-EVA SpO2 94% 91% 92%
Pre-EVA HR 84 91 89
Post-EVA HR 103 120 96
Max Perceived Physical Exertion (1-10) 7 7 5
Max Perceived Mental Exertion (1 (low effort)-10 (max concentration)) 1 3 4
Max Perceived Emotional Distress (1 (no distress)-10 (very distressed)) 1 1 1
HR @ Max Perceived Exertion 164 150 127
Max HR
Sp02 at Max Perceived Exertion 98% 98% 95%

Subjective Improvements for Future EVAs

  • Secure ear-pieces, hair and sun-protection better
  • Pre-EVA blister prevention suggested prior to next EVA
  • Secure sunglasses